If m=5 kg and M=8 kg, find the acceleration of the masses and the tension of the rope


1
Expert's answer
2021-06-24T07:20:12-0400

"8g-T=8a\\
 T-g=5a"

Adding the two equations gives

"3g=13a\\



a=\frac{3g}{13}\\



g=9.8m/s^2\\
a=\frac{3\times 9.8}{13}\\
a=2.26m/s^2\\"

To get the Tension T

"T=5(g+a)\\
T=5(9.8+2.26)\\
T=60.3N"